### v2

@@ Abstract
Results: We developed pangene, a computational tool to identify gene orientation, gene order and gene copy-number changes in a collection of genomes. Pangene aligns a set of input protein sequences to the genomes, resolves redundancies between protein sequences and constructs a gene graph with each genome represented as a walk in the graph. It additionally finds subgraphs that encodes gene content changes. Applied to the human pangenome, pangene identifies known gene-level variations and reveals complex haplotypes that are not well studied before. Pangene also works with high-quality bacterial pangenome and reports similar numbers of core and accessory genes in comparison to existing tools.
@@ 1 Introduction
Thanks to the recent advances in sequencing technologies (Wenger et al., , 2019) and assembly algorithms (Nurk et al., , 2020; Cheng et al., , 2021; Rautiainen et al., , 2023), we can routinely achieve haplotype-resolved assembly over genes under copy-number or order changes. We have also developed algorithms to construct pangenome sequence graphs that encode variations between genomes. However, how to identify these gene-level variations is not straightforward. Among the three pangenome construction tools used by the Human Pangenome Reference Consortium (HPRC), minigraph (Li et al., , 2020) and minigraph-cactus (Hickey et al., , 2023) are unable to align through complex genomic regions and may miss genes in long segmental duplications; PGGB (Garrison et al., , 2023) collapses paralogous genes which makes it difficult to study individual paralogs. In addition, all three tools do not directly reveal how genomic variations affect genes. To study gene-level variations, HPRC had to manually annotate genes on each haplotype (Liao et al., , 2023) which is a time-consuming process. The current human pangenome tooling is not designed for studying gene variations.
@@ 2 Methods
Given perfect gene annotation and orthology assignment between the genes, the pangene graph construction algorithm is conceptually simple: it takes an orthologous group as a vertex and adds an edge between two groups if on an input genome, a gene in one group is adjacent to a gene in the other group (Fig. 1). The practical difficulty is to obtain accurate annotation in the presence of redundant sequences, paralogous genes and errors in assembly or alignment. On the theoretical side, bubble finding is an unsolved problem on its own and will also be a major topic of this article.
@@ 2.1 Defining pangene graphs
Figure 2: An example of a gene graph. (a) A gene graph for visualization. (b) The corresponding GFA format. (c) Bidirected graph. (d) Biedged graph. (e) Directed graph. (f) Net graph, which loses information. (g) A depth-first traversal of the net graph with “0” representing a super node not in the original net graph.
@@
As pangene constructs the graph from protein-to-genome alignment, it also attaches alignment scores to edges. Let $T_{x\to y}\subset T$ be the set of contigs that support $x\to y$ and $S(x|y,t)$, $t\in T_{x\to y}$, be the alignment score of $x$ when $x\to y$ is supported by contig $t$. $S(x|y)$ is calculated as
@@ 2.4 Adjusting pangene graphs
First, a gene may be wrongly mapped to its paralogs if its best position is missing. Suppose a gene on chromosome X has a second best hit on an autosome. If the alignment score of the autosomal hit is lower than the hit to chromosome X by 3% miniprot will only report the hit on chromosome X. However, because a phased male human assembly does not have chromosome X, the gene will be mapped to the autosome (Fig. 3a). This false hit will lead to spurious edges in the initial graph (Fig. 3b). To solve this problem, pangene marks $x\to y$ as a false edge if there exists $x\to z$ such that $y$ and $z$ are on different contigs in every input genome and $S(x|y)<S(x|z)\cdot 95\%$. Pangene then filters out all alignment of $x$ that are incident to a false edge $x\to y$ (Fig. 3c).
@@
Both cases may also be caused by processed pseudogenes. When a protein-coding gene has a recent pseudogene that has a near identical sequence, the two heuristics may not work due to the score ratio threshold. By default, pangene filters out an alignment without splicing if the protein has a spliced alignment in other samples. Pangene also has an option to drop a protein if it does not have spliced alignment in any sample. For human samples, these strategies can reduce false connections between chromosomes.
@@ 2.5 Past work on bubble finding
Weak superbubble is defined on directed graphs, not on bidirected graphs. If we apply the previous algorithms to directed graph $G_{D}$, we will find most superbubbles twice on separate strands. Superbubble cannot capture the visual intuition in the presence of inversions, either. For example, Fig. 4a seems to have a bubble visually but the corresponding directed graph (Fig. 4f) does not have a superbubble. Dabbaghie et al., (2022) implemented the algorithm by Onodera et al., (2013) for sequence graphs, but how the issues above are handled is not apparent. The closest concept to weak superbubble in bidirected graphs is snarl (Paten et al., , 2018), which leads to a bidirected subgraph that can be separated from the rest of the graph. Snarl does not require reachability which is necessary in defining superbubbles. To this end, there are no equivalent definition of weak superbubble in directed graphs.
@@ 3.1 Structural variants between two human genomes
We downloaded GenCode comprehensive human annotation v45, retained the protein-coding transcripts tagged with “Ensembl_canonical” and filtered out readthrough transcripts and mitochondrial genes. Only one transcript was selected per gene. Because GenCode does not include HLA-DRB3, HLA-DRB4 and several KIR genes, we manually added 20 HLA-DRB genes and alleles from the IPD-IMGT/HLA database and 15 KIR genes from the IPD-KIR database. In the end, we constructed a protein set with 19,421 sequences.
@@ 3.2 Analyzing 100 human haplotypes
Figure 6: Human and great ape haplotypes around LRRC37A* genes. GRCh38 has LRRC37A and LRRC37A2 in 17q21.31 and LRRC37A3 in 17q24.1. All great apes have only one LRRC37A* gene. A dashed line links two genes that are close but not immediately adjacent to each other.
We obtained the assembly of CN1 (Yang et al., , 2023), YAO (He et al., , 2023) and 47 samples released by the Human Reference Pangenome Consortium (Liao et al., , 2023). Together with GRCh38 and T2T-CHM13, this gave us 100 human haplotypes. We aligned the same set of proteins to the 100 haplotypes. Pangene took less than one minute to construct a graph and less than one second to identify 266 generalized bibubbles in the graph. Many of the bibubbles were supported by one contig only. Pruning edges supported by one contig resulted in a graph containing 209 generalized bibubbles. Further dropping about 2,000 single-coding-exon genes led to a graph containing 163 bibubbles. These included manually confirmed gene-level variations around CYP2D6/CYP2D7, C4A/C4B and RHD/RHCE (Liao et al., , 2023).
@@
Table 2: Bacterial pangenome analysis
@@
Ecoli: #total genes 13,723 13,009 14,630
Ecoli: #core genes 2,893 3,081 2,966
@@
Pangene constructed a graph consisting of 4,216 genes, 3,652 of which were present in all 152 genomes (Table 2). To check if pangene captured the gene content in these strains, we compared the pangene result to Panaroo v1.3.4 in the strict mode. We aligned the Panaroo proteins to the pangene proteins with MMseqs2 (Steinegger and Söding, , 2017) v13.45111 and identified 46 Panaroo proteins do not hit to pangene proteins. We mapped the 46 proteins to H37Rv with miniprot and found 43 of them can be aligned and 76% of the aligned regions overlap with annotated CDS in RefSeq. Manually investigating the overlaps revealed that most of the 43 proteins were aligned to the opposite strand of some RefSeq genes or in different reading frames. Identifying homology based on the genomic locations of input proteins, pangene did not include them into the final graph. According to the RefSeq annotation, only 0.2% of coding regions in H37Rv are present in more than one genes – overlapping genes are rare in M. tuberculosis.
We additionally ran PPanGGOLiN (Gautreau et al., , 2020) v1.2.105 with the Prokka annotation as the input. PPanGGOLiN collected 4,744 genes in the pangenome with 3,463 present in all. 277 genes did not hit to genes selected by pangene. 274 of these genes could be aligned to H37Rv by miniprot and 90% of the aligned regions overlap with annotated CDS in RefSeq. PPanGGOLiN tends to select more overlapping genes in different reading frames, confirming the observation by Tonkin-Hill et al., (2020).
@@
M. tuberculosis has low diversity with each strain similar to each other (Marin et al., , 2022). To understand how pangene performs given more diverse strains, we collected 50 E. coli genomes with complete assemblies (Shaw et al., , 2021). We did not run Prokka on this dataset but instead used the gene annotation provided by NCBI.
We followed the same procedure to run pangenome tools. To get clean graph, pangene by default filters out genes that had $>$10 edges or connected $>$3 distant loci in the graph. This filter only removed six genes in M. tuberculosis dataset but it filtered out 627 genes in E. coli. We added option “-p.001 -g50 -r10” to retain more genes. Although pangene collected fewer genes (Table 2), only 98 Panaroo genes do not hit to genes collected by pangene. The differences between the tools might be determined by subtle thresholds on how to resolve homology and may not reflect the capability of each algorithm.
@@ 4 Discussions
Pangene constructs a graph to jointly represent the gene content of multiple high-quality genomes. It provides approximate gene annotation and orthology assignment as a byproduct. Given bacterial genomes, it plays a similar role to other bacterial pangenome tools for identifying core and accessory genes (Tonkin-Hill et al., , 2023). Pangene is unique in that it also works for eukaryotic genomes and captures localized gene-level variants. Pangene complements whole-genome alignment based pangenome tools and gives a more concise high-level view of gene content changes, though it ignores small variants and non-allelic gene conversions (Vollger et al., , 2023).
The pangene graph construction algorithm is ad hoc. We have manually tuned various heuristics to correctly encode known gene-level variants in human, but may have overlooked other cases that pangene may fail to represent. Ideally, we would prefer to model graph construction to a global optimization problem. We have not been able to find such a formulation that can reliably encode known variations.
As protein-to-genome alignment can identify orthology within mammals or even vertebrates (Li, , 2023), we could in principle apply pangene to more distantly related species. We did construct a gene graph from the human and mouse reference genomes. The graph revealed synteny blocks visible from the human-mouse whole-genome alignment, but the topology looked complex due to frequent large-scale rearrangements over long evolutionary distance. Many generalized bibubbles identified from this graph were due to real orthologous alignments missed by miniprot. We will need another way at higher level to investigate such a graph. Another practical challenge in building a cross-species graph is to select the set of input proteins. When constructing the great ape graph, we tried to merge human and great ape protein sequences. In the resultant graph, we observed thousands of alignments in the intergenic or intronic regions of GRCh38 which degraded the overall quality of the graph. The issue will become more prominent when we take diverse genomes without high-quality gene annotation as input. Nevertheless, we note that cross-species whole-genome alignment is also challenging and is perhaps no better than pangene. We still believe with improvements, pangene can construct a cross-species graph that helps to reveal evolution in longer term.

### v3

@@ Abstract
Results: We developed pangene, a computational tool to identify gene orientation, gene order and gene copy-number changes in a collection of genomes. Pangene aligns a set of input protein sequences to the genomes, resolves redundancies between protein sequences and constructs a gene graph with each genome represented as a walk in the graph. It additionally finds subgraphs, which we call bibubbles, that capture gene content changes. Applied to the human pangenome, pangene identifies known gene-level variations and reveals complex haplotypes that are not well studied before. Pangene also works with high-quality bacterial pangenome and reports similar numbers of core and accessory genes in comparison to existing tools.
@@ 1 Introduction
Thanks to the recent advances in sequencing technologies (Wenger et al., , 2019) and assembly algorithms (Nurk et al., , 2020; Cheng et al., , 2021; Rautiainen et al., , 2023), we can routinely achieve haplotype-resolved assembly over genes under copy-number or order changes. We have also developed algorithms to construct pangenome sequence graphs that encode variations between genomes. However, how to identify these gene-level variations is not straightforward. Among the three whole-genome pangenome construction tools used by the Human Pangenome Reference Consortium (HPRC), minigraph (Li et al., , 2020) and minigraph-cactus (Hickey et al., , 2023) are unable to align through complex genomic regions and may miss genes in long segmental duplications; PGGB (Garrison et al., , 2023) collapses paralogous genes which makes it difficult to study individual paralogs. In addition, all three tools do not directly reveal how genomic variations affect genes. To study gene-level variations, HPRC had to manually annotate genes on each haplotype (Liao et al., , 2023) which is a time-consuming process. PGR-TK (Chin et al., , 2023) reconstructs local haplotype structures from genomic sequences but it is does not directly model genes and is not intended for whole-genome data. The current human pangenome tooling is not designed for studying gene variations.
@@ 2 Methods
Given perfect gene annotation and orthology assignment between the genes, the pangene graph construction algorithm is conceptually simple: it takes an orthologous group as a vertex and adds an edge between two groups if on an input genome, a gene in one group is adjacent to a gene in the other group (Fig. 1). The practical difficulty is to obtain accurate annotation in the presence of redundant sequences, paralogous genes and errors in assembly or alignment.
An important use of pangenome graphs is to identify genomic variations represented as “bubbles”. While bubbles have been extensively studied in directed graphs (Onodera et al., , 2013), they have not been rigorously defined on bidirected graphs graphs that are the necessary outputs of assemblers and pangenome constructors. If the field is to move towards the more generalized bidirected graph as a representation of genomic variation developing the concept of bubbles in this context will be necessary. Hence, how to define and to find “bubbles” in bidirected graphs will be a major topic of this article.
@@ 2.1 Defining pangene graphs
Figure 2: An example of a gene graph. (a) A gene graph for visualization. (b) The corresponding GFA format. (c) Bidirected graph. (d) Biedged graph. (e) Directed graph. (f) Net graph, which loses information. (g) A depth-first traversal of the net graph with “0” representing a super node not in the original net graph. (a) through (e) are equivalent representations.
@@
As pangene constructs the graph from protein-to-genome alignment (Li, , 2023), it also attaches alignment scores to edges. Let $T_{x\to y}\subset T$ be the set of contigs that support $x\to y$ and $S(x|y,t)$, $t\in T_{x\to y}$, be the alignment score of $x$ when $x\to y$ is supported by contig $t$. $S(x|y)$ is calculated as
@@ 2.4 Adjusting pangene graphs
First, a gene may be wrongly mapped to its paralogs if its best position is missing. Suppose a gene on chromosome X has a second best hit on an autosome. If the alignment score of the autosomal hit is lower than the hit to chromosome X by 3% miniprot will only report the hit on chromosome X. However, because a phased male human assembly does not have chromosome X, the gene will be mapped to the autosome (Fig. 3a). This false hit will lead to spurious edges in the initial graph (Fig. 3b). To solve this problem, pangene marks $x\to y$ as a false edge if there exists $x\to z$ such that $y$ and $z$ are on different contigs in every input genome and $S(x|y)<S(x|z)\cdot r_{1}$ ($r_{1}$ defaults to 95%). Pangene then filters out all alignment of $x$ that are incident to a false edge $x\to y$ (Fig. 3c).
@@
Both cases may also be caused by processed pseudogenes. When a protein-coding gene has a recent pseudogene that has a near identical sequence, the two heuristics may not work due to the score ratio threshold. By default, pangene filters out an alignment without splicing as a potential pseudogene alignment if the protein has a spliced alignment in other samples. Pangene also has an option to drop a protein if it does not have spliced alignment in any sample. For human samples, these strategies can reduce false connections between chromosomes.
@@ 2.5 Past work on bubble finding
Weak superbubble is defined on directed graphs, not on bidirected graphs. If we apply the previous algorithms to directed graph $G_{D}$, we will find most superbubbles twice on opposite strands. More importantly, superbubble cannot capture the visual intuition in the presence of inversions. For example, in Fig. 4a, >A and <C enclose a subgraph that looks like a bubble, but the equivalent directed graph (Fig. 4f) does not have a superbubble when >A and <A are taken as different vertices. Dabbaghie et al., (2022) implemented the algorithm by Onodera et al., (2013) for sequence graphs, but how the issues above are handled is not apparent. The closest concept to weak superbubble in bidirected graphs is snarl (Paten et al., , 2018), which leads to a bidirected subgraph that can be separated from the rest of the graph. Snarl does not require reachability which is necessary in defining superbubbles. To this end, there are no equivalent definition of weak superbubble in directed graphs.
@@ 2.9 Identifying generalized bibubbles
Figure 6: Human and great ape haplotypes around LRRC37A* genes. A red vertical bar indicates the two genes on each side are not adjacent on the genome. In the plot, each genome contains two blocks of genes. The genomic distance between the two blocks is 17.3 Mb in human, 61.3 Mb in bonobo, 17.6 Mb in gorilla or 2.8 Mb in orangutan. GRCh38 has LRRC37A and LRRC37A2 in 17q21.31 and LRRC37A3 in 17q24.1. HG00741 has a known inversion around MAPT. All great apes have only one LRRC37A* gene. Chimpanzee and bonobo have the same gene haplotype. The two orangutan subspecies also have the same haplotype. The LRRC37A3 alignment in gorilla has two in-frame stop codons. It is challenging to distinguish LRRC37A* paralogs as they are similar in sequence and lack synteny with nearby genes.
@@ 3 Results
We applied pangene to HPRC samples and manually inspected subgraphs around CYP2D6, RHD, C4, HLA-DRB1 and KIR genes which have been curated for these samples (Liao et al., , 2023; Zhou et al., , 2024). We also looked at many other genes that are known to have gene-level variations (Sudmant et al., , 2010; Boettger et al., , 2012; Handsaker et al., , 2015). While these genes have not been manually annotated for HPRC samples, we could check whether pangene captured known events such as the MAPT inversion and the ORM1 copy-number changes.
Due to the lack of genome-wide ground truth in human data, we could only manually evaluate pangene in individual subgraphs. To get a better sense of overall statistics, we additionally applied pangene to two bacterial datasets and compared the results to existing bacterial pangenome tools.
@@ 3.1 Structural variants between two human genomes
We downloaded GenCode comprehensive human annotation v45, retained the protein-coding transcripts tagged with “Ensembl_canonical” and filtered out readthrough transcripts (a transcript that joins two adjacent genes) and mitochondrial genes. Only one transcript was selected per gene. Because GenCode does not include HLA-DRB3, HLA-DRB4 and several KIR genes, we manually added 20 HLA-DRB genes and alleles from the IPD-IMGT/HLA database and 15 KIR genes from the IPD-KIR database. In the end, we constructed a protein set with 19,421 sequences.
@@ 3.2 Analyzing 100 human haplotypes
We obtained the assembly of CN1 (Yang et al., , 2023), YAO (He et al., , 2023) and 47 samples released by the Human Reference Pangenome Consortium (Liao et al., , 2023). Together with GRCh38 and T2T-CHM13, this gave us 100 human haplotypes. We aligned the same set of proteins to the 100 haplotypes. Pangene took less than one minute to construct a graph and less than one second to identify 266 generalized bibubbles in the graph. Many of the bibubbles were supported by one contig only. Pruning edges supported by one contig resulted in a graph containing 209 generalized bibubbles. Further dropping about 2,000 single-coding-exon genes led to a graph containing 163 bibubbles. These included manually confirmed gene-level variations around CYP2D6, C4 and RHD (Liao et al., , 2023) and around HLA-DBR1 (Zhou et al., , 2024).
Pangene effectively annotates each input genome. For a sanity check, we compared the pangene gene annotation to the GENCODE annotation. Out of 19,043 protein coding genes in both annotations, only four gene locations are different. In one example, pangene assigned NUDT4B to NUDT4. Pangene misses NUDT4B because it is a single-exon gene, and pangene puts NUDT4B at the NUDT4 locus because the protein sequence of NUDT4 is a strict substring of NUDT4B.
We also compared the pangene annotation to the T2T-CHM13 annotation generated by Ensembl. There are 41 differences out 18,676 genes present in both annotations. These include four genes CSAG2, CSAG3, MAGEA6 and MAGEA3 that are close to each other. Pangene gives better protein-to-genome alignment for the four genes. We speculate that Ensembl may be annotating these genes based on synteny with GRCh38, but synteny here might be broken due to an apparent inversion around the genes. For another example, in the RCCX gene cluster (Carrozza et al., , 2021), pangene annotates CYP21A2 ahead of TNXB but Ensembl puts CYP21A2 inside an intron of TNXB, which seems to be an error. Ensembl also misses C4A, another gene in RCCX. We believe pangene is doing better in these cases. Most other genes among the 41 differences come from huge complex gene clusters. We cannot tell what the correct annotation is.
@@
Our analysis is insensitive to thresholds $r_{1}$ and $r_{2}$ in Section 2.4. Reducing $r_{1}$ from 95% to 90% and $r_{2}$ from 98% to 96% only affects 9 out of 266 bibubbles.
3.3 Incorporating 10 great ape haplotypes
@@
Table 2: Number of detected genes in bacterial pangenomes
@@
Ecoli: #total genes 14,342 13,902 14,291
Ecoli: #core genes 3,065 3,118 3,038
@@
Pangene constructed a graph consisting of 4,216 genes, 3,652 of which were present in all 152 genomes (Table 2; Fig. 7). To check if pangene captured the gene content in these strains, we compared the pangene result to Panaroo v1.3.4. We aligned the Panaroo proteins to the pangene proteins with MMseqs2 (Steinegger and Söding, , 2017) v13.45111 and identified 46 Panaroo proteins do not hit to pangene proteins at the default E-value threshold of $10^{-3}$. We mapped the 46 proteins to H37Rv with miniprot and found 43 of them can be aligned and 76% of the aligned regions overlap with annotated CDS in RefSeq. Manually investigating the overlaps revealed that most of the 43 proteins were aligned to the opposite strand of some RefSeq genes or in different reading frames. Identifying homology based on the genomic locations of input proteins, pangene did not include them into the final graph. Conversely, 40 pangene proteins do not hit to Panaroo proteins.
We additionally ran PPanGGOLiN (Gautreau et al., , 2020) v1.2.105 with the Prokka annotation as the input. PPanGGOLiN collected 4,744 genes in the pangenome with 3,463 present in all. 277 genes did not hit to genes selected by pangene. 274 of these genes could be aligned to H37Rv by miniprot and 90% of the aligned regions overlap with annotated CDS in RefSeq. We note that although overlapping genes in different open reading frames are rare in H37Rv (0.2% of coding regions), these genes may occur in other strains and may have functional implications (Vargas et al., , 2023; Snobre et al., , 2024). PPanGGOLiN’s preference to count such genes in the pangenome is not necessarily a weakness.
@@
M. tuberculosis has low diversity with each strain similar to each other (Marin et al., , 2022). To understand how pangene performs given more diverse strains, we collected 50 E. coli genomes with complete assemblies (Shaw et al., , 2021).
We followed the same procedure to run pangenome tools. To get clean graph, pangene by default filters out genes that had $>$10 edges or connected $>$3 distant loci in the graph. This filter only removed six genes in M. tuberculosis dataset but it filtered out 848 genes in E. coli. We added option “-p.001 -g50 -r10” to retain more high-copy genes and genes connecting multiple loci. Although pangene collected fewer genes (Table 2), only 72 Panaroo genes do not hit to genes collected by pangene and the pangene pangenome size after clustering is not noticeably smaller than the Panaroo and the PPanGGOLiN pangenomes (Fig. 7). The differences between the tools might be determined by subtle thresholds on how to resolve homology and may not reflect the capability of each algorithm.
Figure 7: Comparing three bacterial pangenome construction tools. In each venn diagram, pangene, Panaroo and PPanGGOLiN proteins are merged and clustered with MMseqs2 (easy-cluster -c 0.9 --min-seq-id 0.5). The diagram shows the number of clusters in each intersection. MMseqs2 merges paralogous genes, so the number of clusters for each tool is smaller than the number of genes in Table 2. Venn diagrams were generated with BioVenn (Hulsen et al., , 2008).
@@ 4 Discussions
Pangene constructs a graph to jointly represent the gene content of multiple high-quality genomes. It provides approximate gene annotation and orthology assignment as a byproduct. Given bacterial genomes, it plays a similar role to other bacterial pangenome tools for identifying core and accessory genes (Tonkin-Hill et al., , 2023). Pangene is unique in that it also works for eukaryotic genomes and captures localized gene-level variants. Pangene complements whole-genome alignment based pangenome tools and gives a more concise high-level view of gene content changes.
Constructing a pangene graph is challenging because the biologically correct graph is unknown. We tend to think a “correct” graph should model the evolution history in principle. However, evolution is complex in the presence of convergent/divergent evolution in the long time scale and non-homologous recombinations and non-allelic gene conversions (Vollger et al., , 2023) in the short time scale. We do not know the evolution history of most complex loci except several well-studied cases. Furthermore, we take each gene as an atomic unit but biologically, a gene may be subjected to structural changes such as losing/gaining exons or fusing with another gene. These events are difficult to model even if we had the exact annotation in all input genomes.
The pangene graph construction algorithm is ad hoc. We have manually tuned various heuristics to correctly encode known gene-level variants in human, but may have overlooked other cases that pangene may fail to represent. Ideally, we would prefer to model graph construction to a global optimization problem. We have not been able to find such a formulation that can reliably encode known variations. This limitation is not specific to pangene; it is also applied to other pangenome construction tools.
As protein-to-genome alignment can identify orthology within mammals or even vertebrates (Li, , 2023), we could in principle apply pangene to more distantly related species. We did construct a gene graph from the human and mouse reference genomes. The graph revealed synteny blocks visible from the human-mouse whole-genome alignment, but the topology of the whole graph looked complex due to frequent large-scale rearrangements over long evolutionary distance. Many generalized bibubbles identified from this graph were due to real orthologous alignments missed by miniprot. We will need another way at higher level to investigate such a graph. Another practical challenge in building a cross-species graph is to select the set of input proteins. When constructing the great ape graph, we tried to merge human and great ape protein sequences. In the resultant graph, we observed thousands of alignments in the intergenic or intronic regions of GRCh38 which degraded the overall quality of the graph. The issue will become more prominent when we take diverse genomes without high-quality gene annotation as input. Nevertheless, we note that cross-species whole-genome alignment is also challenging and is perhaps no better than pangene. We still believe with improvements, pangene can construct a cross-species graph that helps to reveal evolution in longer term.